OpenStudy (anonymous):

Can someone help me understand this limit please?: lim sinx/x as x->-infinity

OpenStudy (anonymous):

\[-1<\sin x<1\] \[\frac{ -1 }{ x }<\frac{ \sin x }{ x }<\frac{1}{x}\] squeeze theorem

OpenStudy (anonymous):

\[\lim_{x \rightarrow -\infty}\frac{ 1 }{ x }=0\]

OpenStudy (anonymous):

When we do this, why does it become 0? I understand squeeze theorem, but I am just a bit confused. Do you plug a zero in for x or?

OpenStudy (anonymous):

 what happens asa you increase in the -x direction,what is the y value approaching

OpenStudy (anonymous):

So the x value is going to coming very close to 0 but never reach it? Is the only way to determine this by a graph, or is their some special rule that I am forgetting?

OpenStudy (anonymous):

\[x \to \infty ,y \to 0\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 },\frac{ 1 }{ 8 },\frac{ 1 }{ 100 },\frac{ 1 }{ 10000000000 }\] x is increasing but y is converging to a certain value can you see it

OpenStudy (anonymous):

0